What is the freezing point of solution prepared by disolving 15 g of KBr in 100 ml of water? The freezing point depression constant of water is 1.86.


1
Expert's answer
2020-10-21T13:18:35-0400

ΔTf=Kf×Cm

Kf(H2O)=1.86

Cm (molality) is number of moles of solute per kilogram solvent.

ΔTf=Kf×m

M (KBr) = 119 g/mol

n (NaCl) = 15/119 = 0.13 mol

Density of water is 1 kg/l. Threfore, mass of 100 ml = 0.1 L of water is 0.1 kg.

Сm (KBr) = 0.13/0.1=1.3 mol

ΔTf=1.86×1.3=2.4

Therefore Tf of an aqueous solution containing 15 g of KBr in 100 ml of H2O will be -2.4°C